I have four questions.
1) Find and label the vertex and the line of symetry f(x)=1/3x^2
2) Find the line of symetry f(x)=(x+2)^2
3) Nature of the solutions x^2-6=0 I say 2 real solutions am I correct?
4) Number of tickets sold each day is given by N(x)=-0.3x^2 +7.2x+14 where x is the number of days since the show was announced. When will the daily tickets sales peak? How many tickets will be sold on the first day?
Thanks in advance

Answer by stanbon(75887)

1) Find and label the vertex and the line of symetry f(x)=1/3x^2
x^2 = 3y
(x-0)^2 = 3(y-0)
Vertex is (0,0)
Axis of symmetry is x = 0
==============================================
2) Find the line of symetry f(x)=(x+2)^2
(x+2)^2 = y-0
Since vertex is (-2,0), the axis of symmetry is x = -2
==============================================
3) Nature of the solutions x^2-6=0 I say 2 real solutions am I correct?
x = sqrt(6) and x = -sqrt(6)
Two Real Number solutions.
==============================================
4) Number of tickets sold each day is given by N(x)=-0.3x^2 +7.2x+14 where x is the number of days since the show was announced.
When will the daily tickets sales peak?
max occurs at x = -b/2a = -7.2(2*-0.3) = 12
Peak sales 12 days after the show was announced.
-------------------------------------------------------
How many tickets will be sold on the first day?
N(0) = 14
===============
Cheers,
Stan H.